OpenStudy (anonymous):

What is the quotient: (6x^4 – 15x^3 – 2x^2 – 10x – 4) ÷ (3x^2 + 2) ? can someone walk me through how to solve this step by step?

OpenStudy (therealmeeeee):

(6x^4 - 15x^3 - 2x^2 - 10x - 4) ÷ (3x^2 + 2) = (6x^4 - 15x^3 + 4x^2 - 6x^2 - 10x - 4) ÷ (3x^2 + 2) = (6x^4 + 4x^2 - 15x^3 - 10x - 6x^2 - 4) ÷ (3x^2 + 2) = 2x^2(3x^2 + 2) - 5x(3x^2 + 2) - 2(3x^2 + 2) ÷ (3x^2 + 2) = (3x^2 + 2)(2x^2 - 5x - 2) ÷ (3x^2 + 2) = 2x^2 - 5x - 2 Ans.

OpenStudy (therealmeeeee):

do this help ???